OpenStudy (anonymous):

i need two equivalent fraction for this each frac 1/6, 1/4, 5/9, 1/2, 3/11, 6/12, 11/13, 2/3

OpenStudy (espex):

Is the question that you need to come up with two equivalent fractions for each in the list?

OpenStudy (anonymous):

yes

OpenStudy (espex):

Can you use multiples?

OpenStudy (anonymous):

yes

OpenStudy (espex):

Well the easiest way to do this then would be to pick a number and multiply the numerator and denominator by it., say 1/6 * 2/2 equals 2/12 etc.

OpenStudy (anonymous):

1/6 = 2/12, 3/18 1/4 = 2/4, 3/12 5/9 = 10/18, 15/27 1/2 = 2/4, 3/6 3/11 = 6/22, 9/33 6/12 = 1/2, 12/24 11/13 = 22/26, 33/39 2/3 = 4/6, 6/9

OpenStudy (anonymous):

thankyou!!!!